OpenStudy (anonymous):

Which of the triangles described in the table is a right triangle. A. Q B. R C. S D. T

OpenStudy (anonymous):

Q: 25,20,15 R:26,20,46 S:25,20,1025 T:25,26,650

OpenStudy (alexandervonhumboldt2):

use the pythogorian theorem: a^2+b^2=c^2 c is the hypotenuse (the biggest side). a and b are legs. example: 20^2+15^2=25^2. check is this equation is correct. do this with every triangle.

OpenStudy (alexandervonhumboldt2):



OpenStudy (anonymous):

thanks dude By the way... I'M BATMAN